OpenStudy (anonymous):

If 2 - a, then a -2. True False

OpenStudy (anonymous):

true

OpenStudy (texaschic101):

false sub in any number for "a "....lets pick 5 2 - a = 2 - 5 = -3 a - 2 = 5 - 2 = 3